OpenStudy (anonymous):

You all will never fail to amaze me when u help The functions f(x) and g(x) are defined below. f(x) = -14x + 7 g(x) = -13x + 1 Using a table of values, determine the solution to the equation f(x) = g(x). A. x = 6 B. x = -6 C. x = 8 D. x = -77

OpenStudy (aravindg):

f(x) = -14x + 7 g(x) = -13x + 1 f(x)=g(x) So equate both.

OpenStudy (aravindg):

Solve for x